# for Generative QA

### v1

@@ 3.2 Approach 2: Scaling-$\beta$-binning
Scaling-$\beta$-binning is a standalone posthoc $\beta$-calibration algorithm, that is based on performing a scaling step prior to $\beta$-binning. In the original scaling-binning approach (Kumar et al., 2019), confidence scores are first scaled to their (maximum likelihood) fitted values using logistic regression, and the fitted values are then used as proxy ground truth for histogram binning. Adapting this paradigm, in our scaling-$\beta$-binning, we use a scaling subroutine (defined below) to produce fitted values, which are then used as proxy ground truth for $\beta$-binning (e.g., $t$ in Algorithm 1is set to these fitted values).
@@
We adapt the scaling procedure from Kumar et al. (2019) with a crucial change. The confidence score distributions in different partitions $s\in{\mathcal{S}}$ may have different miscalibration profiles, since an LM may be underconfident for a partition but overconfident for another. While scaling is useful for average-case recalibration (Platt, 1999), for $\beta$-calibration, we propose to use a hierarchical logistic regression model with partial pooling (Goldstein, 2011). Hierarchical model distinguishes between fixed effects (consistent across different partitions) and random effects (allowing for variations at different partitions). The latter explains the variability between partitions that may not be captured by fixed effects alone. Partial pooling allows for information sharing across partitions, which improves estimates for partitions with few data points. In our kd-tree instantiation, as the maximum depth parameter $d$ increases, some partitions may have only a few data points.
We next describe the training and test processes of scaling-$\beta$-binning.
@@
Algorithm 3 Scaling-$\beta$-binning: Train-time Subroutine
@@
We first split the calibration dataset $D$ (by default equally) into 2 sets, $D^{1}$ and $D^{2}$, and invoke Algorithm 3 on $D^{1}$ and $D^{2}$. Similar to the $\beta$-binning approach (Algorithm 1), this produces a hash table ${\mathcal{G}}$ of $g_{\text{UMD}}$’s. The partition of each instance $s$ determines which random intercept $U_{s}$ and slope $V_{s}$ to use. In our kd-tree instantiation, if a point lies outside of bounded kd-tree spaces, we set the random effects to zero, thus assuming that this point behaves similarly to the overall population average.
Test Process. Similar to $\beta$-binning, for a test input $(q_{\text{test}},a_{\text{test}},h_{\text{test}})$, we invoke Algorithm 2.
Due to the scaling step, the computational cost of scaling-$\beta$-binning training process is more expensive than that of $\beta$-binning. Both schemes have the same test process.
3.3 Distribution Free Analysis of $\beta$-binning and Scaling-$\beta$-binning
In the next result, we prove a high probability bound on the $\beta$-calibration error (Definition 2.4) of the above schemes. To formalize the guarantee, we adapt the conditional calibration notion from Gupta et al. (2020); Gupta and Ramdas (2021) to the $\beta$-calibration setting, defined below: [^5] One could define $(\epsilon,\alpha)$-marginal $\beta$-calibration: $\mathbb{P}\left(\left|\mathbb{E}[Y\mid h(Q,A),\beta(Q,A)]-h(Q,A)\right|\leq\epsilon\right)\geq 1-\alpha$. Conditional calibration is a stronger definition than marginal, as it requires the deviation between $\mathbb{E}[Y\mid h(Q,A),\beta(Q,A)]$ and $h(Q,A)$ to be at most $\epsilon$ for every $(s,r)$, including rare ones, not just on average.
Definition 3.1 (Conditional $\beta$-calibration).
Let $\epsilon,\alpha\in(0,1)$ be some given levels of approximation and failure respectively. Confidence function $h:{\mathcal{Q}}\times{\mathcal{A}}\rightarrow[0,1]$ is $(\epsilon,\alpha)$-conditionally $\beta$-calibrated for discretization scheme $\beta:{\mathcal{Q}}\times{\mathcal{A}}\rightarrow S$ if for every distribution $P$ defined in Eq. ( 1),
@@ 3.3 Distribution Free Analysis of $\beta$-binning and Scaling-$\beta$-binning
In practice, $\tilde{y}$ can come from two sources: 1) an LM-constructed ground truth (see ground truth proxy in Table 3), and 2) the fitted values from our S step in scaling-$\beta$-binning. Let $\tilde{D}=\{(q_{i},a_{i},h_{i},\tilde{y}_{i})\}_{i\in[N]}$, denote a (proxy) dataset with samples from $P=P_{Q}\times P_{A|q}\times P_{H|q,a}\times P_{\tilde{Y}|q,a}$. To prove calibration guarantees for our schemes, we rely on the following result.
Theorem 3.1 (Distribution-free $\beta$-calibration guarantee).
Consider an input calibration dataset $\tilde{D}$ defined above with misspecification factor $\nu$ from Definition 3.2. Assume that the $h_{i}$’s are distinct, number of points per bin $b\geq 2$, and number of instances within each partition $n_{s}\geq b$ for every $s\in{\mathcal{S}}$. The calibrator $g_{\text{UMD}}$ retrieved in Line 2 of Algorithm 2, trained using Algorithm 1 with input ${\mathcal{D}}=\tilde{D}$, is $(\epsilon,\alpha)$-conditionally $\beta$-calibrated for any $\alpha\in(0,1)$, with $\epsilon=\sqrt{\frac{\log(2N/b\alpha)}{2(b-1)}}+\nu$.
The proof is in Appendix A.3. The dependence on $\approx 1/\sqrt{b}$ factor comes because the Algorithm 1 delegates at least $b$ points to every bin. We now discuss how Theorem 3.1 is applicable for both $\beta$-binning and scaling-$\beta$-binning with different $\nu$’s.
Applying Theorem 3.1 to $\beta$-binning & Scaling-$\beta$-binning. In our description of $\beta$-binning (Section 3.1), we assumed $t$ is set to the ground truth $y$ (in Algorithm 1), hence, by definition $\nu=0$. Theorem 3.1 can also be used to choose $b$, see the plots for $\nu=0$ in Figure 2.
If the true labels are not available, then we can still use $\beta$-binning say by using an LM to produce proxy ground truth. In this case, the misspecification constant $\nu$ depends on the data generating process of misspecified labels. When an LM is used to produce proxy ground truth, if there is a hold-out set containing the ground truth, then a bound on $\nu$ can be estimated empirically.
In scaling-$\beta$-binning, where $\tilde{y}$ is set to be the fitted values of a hierarchical logistic regression model, the magnitude of misspecification factor $\nu$ depends on the goodness-of-fit of the fitted values. In practice, we can estimate $\nu$ empirically using a hold-out dataset. This estimate can be used to choose $b$. For some levels of $\epsilon$, the same $\epsilon$ as in the case of $\nu=0$ can be attained by setting $b$ to be a higher number. In our kd-tree instantiation, this amounts to using a smaller maximum depth hyperparameter.
@@ 4 Related Work
Calibration for Language Models. Reinforcement learning from human feedback objective may prioritize adherence to user instructions in dialogue over producing well-calibrated predictions. (Kadavath et al., 2022)Lin et al. (2022) introduced the concept of verbalized confidence that prompts LMs to express confidence directly, focusing on fine-tuning, instead of zero-shot verbalized confidence. Mielke et al. (2022) uses an external calibrator for a white-box large language model. Other methods use consistency measures to improve LM calibration (Lyu et al., 2024). Our experimental setup closely relates to recent works in LM confidence elicitation (Tian et al., 2023; Xiong et al., 2024). These methods lack novel post-hoc calibrators and do not offer the rigorous calibration guarantees that ours provide. Calibration has been shown to impact selective QA performance Kamath et al. (2020), but they focus on uncertainty quantification and assumes that the LM allows access to the model likelihood.
@@
Other Metrics for Measuring Calibration Error. Brier score (Brier, 1950) measures the accuracy of probabilistic predictions but while it can be decomposed into calibration and refinement (Blattenberger and Lad, 1985), it doesn’t directly assess calibration. As a result, a model with lower squared error may still be less well-calibrated. Maximum Calibration Error examines the maximum miscalibration across confidence bins (Guo et al., 2017), but in a QA setting, it faces the same issues as calibration error ($\text{CE}(h)$), as shown in Example 2.1. Through $\beta$-calibration, we present a principled and interpretable calibration target for QA settings.
@@ 5 Experiments
Datasets, Models, and Prompts. We use 5 QA datasets: TriviaQA (Joshi et al., 2017), SciQ (Welbl et al., 2017), BigBench (Srivastava et al., 2022), OpenBookQA (Mihaylov et al., 2018), and MMLU (Hendrycks et al., 2021) (see Table 4 for more details). We use two performant models: Mistral (Jiang et al., 2023) and Gemma (Team et al., 2024). To elicit confidence scores, we use two prompt techniques recently suggested in literature: Verb1S-Top1 & Ling1S-Top1 from Tian et al. (2023). See Table 5, (Appendix B.1) for details about the prompts.
@@
Central to the implementation of posthoc calibration and evaluation of calibration is the availability of a label for a question-and-answer pair — specifically, whether the answer provided by the LM is accurate for the question posed by the user. In practice, a common idea to generate this label is to take an LM provided answer, and then use another LM to assess whether the proposed answer is semantically equivalent to the true (ground truth) answer (Tian et al., 2023). To construct the ground truth proxy for $y$, we use Llama 3.1 (Dubey et al., 2024).
@@
Our Methods. We compare the performance of our calibrators: $\beta$-binning (BB) from Subsection 3.1 and hierarchical scaling-$\beta$-binning (HS-BB) from Subsection 3.2. We also include a fully pooled version of scaling-$\beta$-binning (S-BB), by setting $s[i]$ to a constant (thus, one partition) in Eq. ( 5). To set the hyperparameter minimum number of points per bin $b$ (Algorithm 1), we set an $\epsilon$ that is not too large as per Figure 2 and use root finding with the $\epsilon$ expression in Theorem 3.1 to choose $b$. We then search over a range of $b$’s by allowing for a misspecification range between 00 and $0.05$ and a range of maximum kd-tree depths depending on the size of the dataset such that each partition admits a 3–10 bins. To set $B$ in UMD, we follow the guidelines in Gupta and Ramdas (2021). Note that their bound does not involve misspecification factor.
@@
Baselines. We consider the following baselines: no recalibration (None) which returns state-of-the-art elicited confidence scores (Tian et al., 2023), histogram binning (UMD) (Gupta and Ramdas, 2021), Platt scaling (S) (Platt, 1999), and scaling-binning (S-B) (Kumar et al., 2019). These baselines consist of the state-of-the-art ideas in posthoc calibration. Note that the techniques UMD, S, and S-B, aim to minimize the expected calibration error CE (Definition 2.2) and do not take the partitions induced by $\beta$ into account.
Metrics. In this section, we primarily use two metrics for comparison. First, is the $\beta$-calibration error $\text{CE}(h;\beta)$ (Definition 2.4), which is the metric that our methods (presented in Section 3) optimize for. A lower $\text{CE}(h;\beta)$ indicates a more effective scheme for achieving $\beta$-calibration. As explained in Section 2, $\beta$-calibration error generalizes the expected calibration error ($\text{CE}(h)$).
@@
Second, to measure the downstream impact of using $\beta$-calibration as the confidence calibration notion in a QA setting, we adapt selective question answering (Kamath et al., 2020) to our setting: given a threshold $\gamma\in[0,1]$, the answer $a$ is returned for a question $q$ to the end user if $h(q,a)\geq\tau$, and abstains (no answer is returned) otherwise. [^6] For example, when $\tau=0$, all answers are returned irrespective of the confidence score. We adapt the area under risk-coverage curve, which is a standard way to evaluate selective prediction methods (El-Yaniv and Wiener, 2010) to our setting. Our post-hoc schemes produce discretized confidence scores $h$, often with large number of ties, which lead to unreliable risk-coverage calculations. Therefore, we consider the area under accuracy-confidence curve (AUAC): set grid points between 0 and 1, and for each grid point, record the accuracy (again, based on a ground truth proxy of $y$) of all points with a confidence score greater than or equal to that grid point. This accuracy is then used as the height of the curve. A higher AUAC indicates a more effective scheme for selective QA.
@@
Results. Table 2 shows the performance of the posthoc calibrators on MMLU and BigBench datasets. More results are provided in Tables 6 and 7. Our methods (BB, HS-BB, and S-BB) generally achieve the best $\beta$-calibration error $\text{CE}(h;\beta)$ and best area under the accuracy-confidence curve (AUAC). While the first result, in itself, may not be surprising as our proposed schemes aim to minimize $\text{CE}(h;\beta)$, the gap between our techniques and baselines is significantly huge. For example, notice the difference in the calibration error when just using the SOTA confidence elicitation prompts (None) from (Tian et al., 2023) vs. our schemes in Table 2. Among our schemes, HS-BB generally performs best. This is because a parametric model (especially a partially pooled model like hierarchical scaling) helps reduce the variance of the downstream binning averages.
For selective QA, we again notice that our proposed schemes consistently outperforms the baselines. In some cases, the underlying LLM using confidence elicitation prompts (None) is performant in selective answering and attains high AUAC (similar results using different LLMs were noted by Tian et al. (2023)), but are not well-calibrated as demonstrated by their high $\beta$-calibration score. Our schemes, S-BB and HS-BB are generally the top-two performing schemes for this task with comparable and in many cases better AUAC scores than the None scheme. Since the accuracy-confidence curve is generated by examining the accuracy of answers above a confidence threshold, the results demonstrate the desirable quality that the confidence scores provided by our proposed schemes are better at ranking accurate answers higher than inaccurate ones. Crucially, the lower performance of other baselines (like S-B, S, and B) demonstrates that the advantage of having a better calibration target that comes through our definition of $\beta$-calibration. In particular, our $\beta$-calibration framework identifies the optimal kd-tree depth, which is never equal to zero (the depth corresponding to standard average-case calibration) in our experiments.

### v2

@@ 3.2 Approach 2: Scaling QA binning
Scaling QA binning is a standalone posthoc QA-calibration algorithm, that is based on performing a scaling step prior to QA binning. In the original scaling-binning approach (Kumar et al., 2019), confidence scores are first scaled to their (maximum likelihood) fitted values using logistic regression, and the fitted values are then used as proxy ground truth for histogram binning. Adapting this paradigm, in our scaling QA binning, we use a scaling subroutine (defined below) to produce fitted values, which are then used as proxy ground truth for QA binning (e.g., $t$ in Algorithm 1 is set to these fitted values).
@@
We adapt the scaling procedure from Kumar et al. (2019) with a crucial change. The confidence score distributions in different partitions $s\in{\mathcal{S}}$ may have different miscalibration profiles, since an LM may be underconfident for a partition but overconfident for another. While scaling is useful for average-case recalibration (Platt, 1999), for QA-calibration, we propose to use a hierarchical logistic regression model with partial pooling (Goldstein, 2011). Hierarchical model distinguishes between fixed effects (consistent across different partitions) and random effects (allowing for variations at different partitions). The latter explains the variability between partitions that may not be captured by fixed effects alone. Partial pooling allows for information sharing across partitions, which improves estimates for partitions with few data points. In our kd-tree instantiation, as the maximum depth parameter $d$ increases, some partitions may have only a few data points.
We next describe the training and test processes of scaling QA binning.
@@
Algorithm 3 Scaling QA binning: Train-time Subroutine
@@
We first split the calibration dataset $D$ (by default equally) into 2 sets, $D^{1}$ and $D^{2}$, and invoke Algorithm 3 on $D^{1}$ and $D^{2}$. Similar to the QA binning approach (Algorithm 1), this produces a hash table ${\mathcal{G}}$ of $g_{\text{UMD}}$’s. The partition of each instance $s$ determines which random intercept $U_{s}$ and slope $V_{s}$ to use. In our kd-tree instantiation, if a point lies outside of bounded kd-tree spaces, we set the random effects to zero, thus assuming that this point behaves similarly to the overall population average.
Test Process. Similar to QA binning, for a test input $(q_{\text{test}},a_{\text{test}},h_{\text{test}})$, we invoke Algorithm 2.
Due to the scaling step, the computational cost of scaling QA binning training process is more expensive than that of QA binning. Both schemes have the same test process.
3.3 Distribution Free Analysis of QA binning and Scaling QA binning
In the next result, we prove a high probability bound on the QA-calibration error (Definition 2.4) of the above schemes. To formalize the guarantee, we adapt the conditional calibration notion from Gupta et al. (2020); Gupta & Ramdas (2021) to the QA-calibration setting, defined below: [^5] One could define $(\epsilon,\alpha)$-marginal QA-calibration: $\mathbb{P}\left(\left|\mathbb{E}[Y\mid h(Q,A),\beta(Q,A)]-h(Q,A)\right|\leq\epsilon\right)\geq 1-\alpha$. Conditional calibration is a stronger definition than marginal, as it requires the deviation between $\mathbb{E}[Y\mid h(Q,A),\beta(Q,A)]$ and $h(Q,A)$ to be at most $\epsilon$ for every $(s,r)$, including rare ones, not just on average.
Definition 3.1 (Conditional QA-calibration).
Let $\epsilon,\alpha\in(0,1)$ be some given levels of approximation and failure respectively. Confidence function $h:{\mathcal{Q}}\times{\mathcal{A}}\rightarrow[0,1]$ is $(\epsilon,\alpha)$-conditionally QA calibrated for discretization scheme $\beta:{\mathcal{Q}}\times{\mathcal{A}}\rightarrow S$ if for every distribution $P$ defined in Eq. ( 1),
@@ 3.3 Distribution Free Analysis of QA binning and Scaling QA binning
In practice, $\tilde{y}$ can come from two sources: 1) an LM-constructed ground truth (see ground truth proxy in Table 3), and 2) the fitted values from our S step in scaling QA binning. Let $\tilde{D}=\{(q_{i},a_{i},h_{i},\tilde{y}_{i})\}_{i\in[N]}$, denote a (proxy) dataset with samples from $P=P_{Q}\times P_{A|q}\times P_{H|q,a}\times P_{\tilde{Y}|q,a}$. To prove calibration guarantees for our schemes, we rely on the following result.
Theorem 3.1 (Distribution-free QA-calibration Guarantee).
Consider an input calibration dataset $\tilde{D}$ defined above with misspecification factor $\nu$ from Definition 3.2. Assume that the $h_{i}$’s are distinct, number of points per bin $b\geq 2$, and number of instances within each partition $n_{s}\geq b$ for every $s\in{\mathcal{S}}$. The calibrator $g_{\text{UMD}}$ retrieved in Line 2 of Algorithm 2, trained using Algorithm 1 with input ${\mathcal{D}}=\tilde{D}$, is $(\epsilon,\alpha)$-conditionally QA calibrated for any $\alpha\in(0,1)$, with $\epsilon=\sqrt{\frac{\log(2N/b\alpha)}{2(b-1)}}+\nu$.
The proof is in Appendix A.3. The dependence on $\approx 1/\sqrt{b}$ factor comes because the Algorithm 1 delegates at least $b$ points to every bin. We now discuss how Theorem 3.1 is applicable for both QA binning and scaling QA binning with different $\nu$’s.
Applying Theorem 3.1 to QA Binning & Scaling QA Binning. In our description of QA binning (Section 3.1), we assumed $t$ is set to the ground truth $y$ (in Algorithm 1), hence, by definition $\nu=0$. Theorem 3.1 can also be used to choose $b$, see the plots for $\nu=0$ in Figure 2.
If the true labels are not available, then we can still use QA binning say by using an LM to produce proxy ground truth. In this case, the misspecification constant $\nu$ depends on the data generating process of misspecified labels. When an LM is used to produce proxy ground truth, if there is a hold-out set containing the ground truth, then a bound on $\nu$ can be estimated empirically.
In scaling QA binning, where $\tilde{y}$ is set to be the fitted values of a hierarchical logistic regression model, the magnitude of misspecification factor $\nu$ depends on the goodness-of-fit of the fitted values. In practice, we can estimate $\nu$ empirically using a hold-out dataset. This estimate can be used to choose $b$. For some levels of $\epsilon$, the same $\epsilon$ as in the case of $\nu=0$ can be attained by setting $b$ to be a higher number. In our kd-tree instantiation, this amounts to using a smaller maximum depth hyperparameter.
[…]
@@ 4 Related Work
Calibration for Language Models. Reinforcement learning from human feedback objective may prioritize adherence to user instructions in dialogue over producing well-calibrated predictions. (Kadavath et al., 2022)Lin et al. (2022) introduced the concept of verbalized confidence that prompts LMs to express confidence directly, focusing on fine-tuning, instead of zero-shot verbalized confidence. Mielke et al. (2022) uses an external calibrator for a white-box large language model. Other methods use consistency measures to improve LM calibration (Lyu et al., 2024). Our experimental setup closely relates to recent works in LM confidence elicitation (Tian et al., 2023; Xiong et al., 2024). These methods lack novel posthoc calibrators and do not offer the rigorous calibration guarantees that ours provide. Calibration has been shown to impact selective QA performance Kamath et al. (2020), but they focus on uncertainty quantification and assumes that the LM allows access to the model likelihood.
@@
Other Metrics for Measuring Calibration Error. Brier score (Brier, 1950) measures the accuracy of probabilistic predictions but while it can be decomposed into calibration and refinement (Blattenberger & Lad, 1985), it doesn’t directly assess calibration. As a result, a model with lower squared error may still be less well-calibrated. Maximum Calibration Error (MCE) examines the maximum miscalibration across confidence bins (Guo et al., 2017), but in a QA setting, it faces the same issues as calibration error ($\text{CE}(h)$), as shown in Example 2.1. We generalize MCE to the QA-calibration setting in Appendix D. Through QA-calibration, we present a principled and interpretable calibration target for QA settings.
@@ 5 Experiments
Datasets, Models, and Prompts. We use 5 QA datasets: TriviaQA (Joshi et al., 2017), SciQ (Welbl et al., 2017), BigBench (Srivastava et al., 2022), OpenBookQA (Mihaylov et al., 2018), and MMLU (Hendrycks et al., 2021) (see Table 4 for more details). We use two performant models: Mistral (Jiang et al., 2023) and Gemma (Team et al., 2024). To elicit confidence scores, we use two prompt techniques recently suggested in literature: Verb1S-Top1 & Ling1S-Top1 from Tian et al. (2023). See Table 5, (Appendix B.1) for details about the prompts. Central to the implementation of posthoc calibration and evaluation of calibration is the availability of a label for a question-and-answer pair — specifically, whether the answer provided by the LM is accurate for the question posed by the user. In practice, a common idea to generate this label is to take an LM provided answer, and then use another LM to assess whether the proposed answer is semantically equivalent to the true (ground truth) answer (Tian et al., 2023). To construct the ground truth proxy for $y$, we use Llama 3.1 (Dubey et al., 2024). See Appendix E for qualitative examples of question-and-answer pair kd-tree partition assignments.
@@
Our Methods. We compare the performance of our calibrators: QA binning (QAB) from Subsection 3.1 and hierarchical scaling QA binning (HS-QAB) from Subsection 3.2. We also include a fully pooled version of scaling QA binning (S-QAB), by setting $s[i]$ to a constant (thus, one partition) in Eq. ( 5). To set the hyperparameter minimum number of points per bin $b$ (Algorithm 1), we set an $\epsilon$ that is not too large as per Figure 2 and use root finding with the $\epsilon$ expression in Theorem 3.1 to choose $b$. We then search over a range of $b$’s by allowing for a misspecification range between 00 and $0.05$ and a range of maximum kd-tree depths depending on the size of the dataset such that each partition admits a 3–10 bins. To set $B$ in UMD, we follow the guidelines in Gupta & Ramdas (2021). Note that their bound does not involve misspecification factor.
@@
Baselines. We consider the following baselines: no recalibration (None) which returns state-of-the-art elicited confidence scores (Tian et al., 2023), histogram binning (UMD) (Gupta & Ramdas, 2021), Platt scaling (S) (Platt, 1999), and scaling-binning (S-B) (Kumar et al., 2019). These baselines consist of the state-of-the-art ideas in posthoc calibration. Note that the techniques UMD, S, and S-B, aim to minimize the expected calibration error CE (Definition 2.2) and do not take the partitions induced by $\beta$ into account.
@@
Metrics. In this section, we primarily use two metrics for comparison. First, is the QA-calibration error $\text{CE}(h;\beta)$ (Definition 2.4), which is the metric that our methods (presented in Section 3) optimize for. A lower $\text{CE}(h;\beta)$ indicates a more effective scheme for achieving QA-calibration. As explained in Section 2, QA-calibration error generalizes the expected calibration error ($\text{CE}(h)$). Second, to measure the downstream impact of using QA-calibration as the confidence calibration notion in a QA setting, we adapt selective question answering (Kamath et al., 2020) to our setting: given a threshold $\gamma\in[0,1]$, the answer $a$ is returned for a question $q$ to the end user if $h(q,a)\geq\tau$, and abstains (no answer is returned) otherwise. [^6] For example, when $\tau=0$, all answers are returned irrespective of the confidence score. We adapt the area under risk-coverage curve, which is a standard way to evaluate selective prediction methods (El-Yaniv & Wiener, 2010) to our setting. Our posthoc schemes produce discretized confidence scores $h$, often with large number of ties, which lead to unreliable risk-coverage calculations. Therefore, we measure the area under the accuracy-confidence curve (AUAC) by sampling grid points from 0 to 1, computing the accuracy (based on ground truth $y$) of points with confidence greater than each grid point. A higher AUAC indicates a more effective selective QA scheme.
Dataset Prompt LLM Calibrator $\text{CE}(h;\beta)$ AUAC MMLU Ling1s-Top1 Mistral QAB (ours) $0.171\pm 0.005$ $0.20\pm 0.023$ MMLU Ling1s-Top1 Mistral HS-QAB (ours) $\mathbf{0.16\pm 0.005}$ $\mathbf{0.269\pm 0.022}$ MMLU Ling1s-Top1 Mistral S-QAB (ours) $0.163\pm 0.003$ $0.19\pm 0.045$ MMLU Ling1s-Top1 Mistral S-B $0.393\pm 0.004$ $0.141\pm 0.003$ MMLU Ling1s-Top1 Mistral S $0.249\pm 0.005$ $0.122\pm 0.002$ MMLU Ling1s-Top1 Mistral B $0.392\pm 0.003$ $0.139\pm 0.007$ MMLU Ling1s-Top1 Mistral None $0.532\pm 0.008$ $\mathbf{0.269\pm 0.007}$ MMLU Ling1s-Top1 Gemma QAB (ours) $0.232\pm 0.005$ $0.211\pm 0.032$ MMLU Ling1s-Top1 Gemma HS-QAB (ours) $0.182\pm 0.005$ $0.26\pm 0.01$ MMLU Ling1s-Top1 Gemma S-QAB (ours) $\mathbf{0.181\pm 0.009}$ $\mathbf{0.275\pm 0.029}$ MMLU Ling1s-Top1 Gemma S-B $0.382\pm 0.006$ $0.184\pm 0.002$ MMLU Ling1s-Top1 Gemma S $0.201\pm 0.009$ $0.19\pm 0.003$ MMLU Ling1s-Top1 Gemma B $0.385\pm 0.003$ $0.183\pm 0.008$ MMLU Ling1s-Top1 Gemma None $0.603\pm 0.006$ $0.249\pm 0.01$ MMLU Verb1s-Top1 Mistral QAB (ours) $0.197\pm 0.004$ $0.198\pm 0.046$ MMLU Verb1s-Top1 Mistral HS-QAB (ours) $\mathbf{0.149\pm 0.004}$ $\mathbf{0.306\pm 0.035}$ MMLU Verb1s-Top1 Mistral S-QAB (ours) $0.16\pm 0.004$ $0.217\pm 0.057$ MMLU Verb1s-Top1 Mistral S-B $0.362\pm 0.005$ $0.139\pm 0.004$ MMLU Verb1s-Top1 Mistral S $0.258\pm 0.005$ $0.126\pm 0.001$ MMLU Verb1s-Top1 Mistral B $0.352\pm 0.004$ $0.132\pm 0.011$ MMLU Verb1s-Top1 Mistral None $0.639\pm 0.006$ $0.255\pm 0.008$ MMLU Verb1s-Top1 Gemma QAB (ours) $0.197\pm 0.006$ $\mathbf{0.345\pm 0.035}$ MMLU Verb1s-Top1 Gemma HS-QAB (ours) $\mathbf{0.151\pm 0.006}$ $0.3\pm 0.035$ MMLU Verb1s-Top1 Gemma S-QAB (ours) $0.161\pm 0.006$ $0.227\pm 0.074$ MMLU Verb1s-Top1 Gemma S-B $0.361\pm 0.007$ $0.133\pm 0.005$ MMLU Verb1s-Top1 Gemma S $0.228\pm 0.005$ $0.137\pm 0.004$ MMLU Verb1s-Top1 Gemma B $0.352\pm 0.004$ $0.144\pm 0.011$ MMLU Verb1s-Top1 Gemma None $0.636\pm 0.007$ $0.258\pm 0.006$
Table 2: Performance of our QA-calibration methods, QA binning (QAB), Pooled scaling QA binning (S-QAB) and Hierarchical scaling QA binning (HS-QAB)), compared to the baselines, UMD, (B,  (Gupta & Ramdas, 2021)), Platt Scaling (S, (Platt, 1999)), and Scaling-binning (S-B, (Kumar et al., 2019)), and prompt-based approach None (Tian et al., 2023) . The bold entry in the columns for $\text{CE}(h;\beta)$ and AUAC represent the best performing schemes for that metric. However, note that in some case the confidence intervals overlap, especially between our schemes.
@@
Results. Table 2 shows the performance of the posthoc calibrators on MMLU and BigBench datasets. More results are provided in Tables 6 and 7. Our methods (QAB, HS-QAB, and S-QAB) generally achieve the best QA-calibration error $\text{CE}(h;\beta)$ and best area under the accuracy-confidence curve (AUAC). While the first result, in itself, may not be surprising as our proposed schemes aim to minimize $\text{CE}(h;\beta)$, the gap between our techniques and baselines is significantly huge. For example, notice the difference in the calibration error when just using the SOTA confidence elicitation prompts (None) from (Tian et al., 2023) vs. our schemes in Table 2. Among our schemes, HS-QAB generally performs best. This is because a parametric model (especially a partially pooled model like hierarchical scaling) helps reduce the variance of the downstream binning averages.
For selective QA, we again notice that our proposed schemes consistently outperforms the baselines. In some cases, the underlying LLM using confidence elicitation prompts (None) is performant in selective answering and attains high AUAC (similar results using different LLMs were noted by Tian et al. (2023)), but are not well-calibrated as demonstrated by their high QA-calibration score. Our schemes, S-QAB and HS-QAB are generally the top-two performing schemes for this task with comparable and in many cases better AUAC scores than the None scheme. Since the accuracy-confidence curve is generated by examining the accuracy of answers above a confidence threshold, the results demonstrate the desirable quality that the confidence scores provided by our proposed schemes are better at ranking accurate answers higher than inaccurate ones. Crucially, the lower performance of other baselines (S-B, S, and B) demonstrates the advantage of QA-calibration. In particular, our QA-calibration framework identifies the optimal kd-tree depth, which empirically is never equal to zero (standard average-case calibration). We observe similar results in Appendix F when $\beta$ is varied. Lastly, we analyze the shift in confidence scores after posthoc calibration in Appendix G and per-partition reliability in Appendix H.